OpenStudy (anonymous):

How do you simplify this expression? 9 • (7 –4)2 + 9

OpenStudy (anonymous):

Are you familiar with the PEMDAS rule.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Here we have to use that.

OpenStudy (anonymous):

9 • (7 –4)2 + 9 9 • (3)2 + 9 9 • 9 + 9 81 + 9 90

OpenStudy (anonymous):

So 90 is the final answer.. Hope this helped you

OpenStudy (anonymous):

Do you have any questions regarding this problem?

OpenStudy (anonymous):

could you check some of my answers to make sure there right.

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

–5 + z = –12? = -7

OpenStudy (anonymous):

Perfect

OpenStudy (anonymous):

8 • g = –32? g=-4

OpenStudy (anonymous):

Excellent :)

OpenStudy (anonymous):

What is the value of –3 • (7 + (–3)2) ÷ 3? = -16

OpenStudy (anonymous):

is it positive 16 or -16 not sure

OpenStudy (anonymous):

So -1 is the answer

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yeah kind of... is this right What is the solution to –55 + q = 7? = -62

OpenStudy (anonymous):

Hey wait. it is positive 62

OpenStudy (anonymous):

–55 + q = 7 q= 7 + 55 q = 62

OpenStudy (anonymous):

wait the other question is not right What is the value of –3 • (7 + (–3)2) ÷ 3? a)13 b)–2 c)–16 d)–48

OpenStudy (anonymous):

Please recheck your question. Is that (-3)^2 or (-3)2?

OpenStudy (anonymous):

sry it's (-3)^2

OpenStudy (anonymous):

Let me show you that

OpenStudy (anonymous):

–3 • (7 + (–3)^2) ÷ 3